Workshop supervisor

F. Frans is the workshop supervisor of NBW. She is paid a fixed salary monthly of N$10000. To keep her motivated she is entitled to an annually bonus of N$9000. NBW contributes N$9000 to her pension fund annually. She works 9 hours per day from Monday to Friday every week. She is also entitled to 28 days paid vacation leave. There are 12 public holidays in the year for which 9 fall on weekends. Assume a 365 days year.

Required:

Using information relating to F. Frans compute Frans s hourly tariff per hour


1
Expert's answer
2022-03-18T12:00:08-0400

Monthly salary = 10000

Annual Bonus = 9000

Pension fund per annum = 9000

working hours per day = 9

vacation leave = 28 days

public holidays = 12

weekend holidays = 9

yearly salary = 10000* 12= 120000

total salary = 120000+ 9000= 129000

working hours per week = 9*5=45

public holiday hours 12-9 =3 *9 = 18

45-18 = 27

number of working hours per year = 27*52 = 1404

the hourly tariff = "= \frac{129000}{1404}=91.88034"

F. Frans earns an approximate of N91.88